# EngChain: A Symbolic Benchmark for Verifiable Multi-Step Reasoning in Engineering

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly being applied to specialized, high-stakes domains like engineering, which demands rigorous evaluation of their complex reasoning capabilities. While current benchmarks assess language understanding, factual recall, mathematics or code generation, none capture the integrative reasoning central to engineering where scientific principles, quantitative modeling and practical constraints must converge. To address this gap, we introduce EngChain, a benchmark for verifiable multi-step engineering problem-solving. EngChain contains 90 problems spanning three engineering branches, organized into 9 domains and 20 distinct areas. The problems are generated from symbolic templates with a high degree of randomization to ensure diversity and eliminate the risk of contamination. With this benchmark, we move beyond final answer accuracy with a two-stage evaluation: we first quantitatively verify the numerical and semantic validity of each reasoning step and then introduce LLM-as-a-Judge, an automated system to qualitatively categorize the identified reasoning errors.
EngChain: A Symbolic Benchmark for Verifiable Multi-Step Reasoning in Engineering
Ayesha Gull1 Muhammad Usman Safder1 Rania Elbadry2 Preslav Nakov2 Zhuohan Xie2 1Namal University 2MBZUAI ProjectCodeApp
**footnotetext: Equal contribution.
@@ 1 Introduction
As large language models (LLMs) transition from general-purpose language tasks to specialized, high-stakes domains such as scientific discovery, industrial design, and engineering analysis, the need for rigorous and meaningful evaluation has become significant (Zhao et al., 2023; Xie et al., 2023a). The progress of the field is gated by the quality of its benchmarks: without tests that capture practical, physically grounded reasoning, it is difficult to separate genuine capability from sophisticated mimicry.
To meet this challenge, we introduce EngChain, a new benchmark for verifiable, multi-step engineering problem-solving. Our benchmark is built on a foundation of symbolic templates, each capable of programmatically generating a vast number of unique, verifiable problems. Figure 1 provides a concrete example from our benchmark, illustrating how a single template generates a complete problem instance along with its full chain-of-thought solution.
@@
Figure 1: An example from EngChain (CSTR Volume Calculation). A symbolic template (1) generates a unique problem instance (2) and its verifiable, step-by-step chain-of-thought solution (3).
This design directly targets two key problems in current LLM evaluation. The first is “benchmark saturation” (Ott et al., 2022), where models rapidly achieve superhuman performance on static benchmarks like GLUE (Wang et al., 2019) and even their more complex successors such as BBH (Kazemi et al., 2025). The second is a structural limitation: most benchmarks evaluate critical skills in disciplinary silos. General knowledge benchmarks (e.g., MMLU (Hendrycks et al., 2020)) test broad factual recall, while specialized benchmarks (e.g., MATH (Hendrycks et al., 2021b) or HumanEval (Chen et al., 2021)) assess abstract logic and algorithmic translation, respectively.
This fragmentation is limiting for engineering which is fundamentally an integrative discipline that requires the synthesis of these isolated skills, combining scientific principles, mathematical modeling, and practical constraints (e.g., Moaveni, 2019; Dym et al., 2012). This holistic, procedural reasoning is not captured by existing benchmarks (e.g., Felten et al., 2025; Li et al., 2025b), which predominantly focus on final-answer accuracy. In engineering, where a flawed process can lead to catastrophic failure, this evaluation gap is critical. To address this gap, EngChain uses symbolic templates to produce integrative reasoning problems and a two-stage evaluation via LLM-as-a-Judge that verifies the full reasoning chain rather than just the final answer.
Verifiable multi-step reasoning is fundamental to building trustworthy AI systems across all domains. Concurrent with our work, FinChain (Xie et al., 2025) introduced a symbolic benchmark for verifiable chain-of-thought financial reasoning. However, the two benchmarks address fundamentally different reasoning types. FinChain evaluates reasoning based on economic principles, whereas EngChain evaluates reasoning grounded in the physical laws and practical constraints of engineering. Moreover, FinChain uses a global alignment metric to score the entire reasoning path, while EngChain employs a two-stage evaluation that provides both a quantitative Step F1-Score for procedural correctness and a qualitative, diagnostic error analysis using our LLM-as-a-Judge system.
In summary, our contributions are threefold:
  1. 1.
    We introduce EngChain, a new symbolic benchmark for verifiable, multi-step engineering problem-solving with 90 problem templates across nine domains.
  2. 2.
    We propose a rigorous two-stage evaluation framework, including our qualitative LLM-as-a-Judge system, that moves beyond final answer accuracy to diagnose reasoning failures.
  3. 3.
    We present a comprehensive analysis of 11 frontier LLMs, revealing a critical “right for the wrong reasons" gap and identifying that the majority of reasoning failures are due to Conceptual Errors, not calculation slips.
@@ 2 Related Work
2.1 Benchmarks in Mathematics and Coding
A large body of work on multi-step reasoning benchmarks has focused on foundational, yet isolated, cognitive skills, particularly in mathematics and coding. In mathematical reasoning, benchmarks ranging from grade-school problems (GSM8K)  (Cobbe et al., 2021) to high-school and college-level competition problems (MATH)  (Hendrycks et al., 2021b) establish a clear standard for evaluating abstract logical deduction. This is further refined by HARDMATH  (Fan et al., 2024) for applied mathematics and GSM-Symbolic  (Mirzadeh et al., 2024) for programming and symbolic tasks. Similarly, benchmarks for coding and software engineering evaluate models across different levels of complexity. HumanEval  (Chen et al., 2021) focuses on translating natural language instructions into small, correct code snippets. MBPP  (Austin et al., 2021) and its derivative MBUPP extend this by testing models on a broader set of programming problems with varying difficulty. SWE-Bench  (Jimenez et al., 2024) addresses complex, real-world codebases, measuring a model’s ability to solve practical software engineering challenges. While these benchmarks are essential for capturing foundational skills, their focus on isolated abilities limits their usefulness for engineering. A model can excel at solving abstract equations or writing pure algorithms yet still be unable to apply those skills to a problem constrained by physical laws and material properties.
2.2 Benchmarks in the Physical Sciences
Recognizing this limitation, later benchmarks focus on applied reasoning within the physical sciences. UGPhysics  (Xu et al., 2025) provides broad coverage and tests how models apply mathematical reasoning to physical laws. Building on this, PhysReason  (Zhang et al., 2025a) emphasizes process-oriented problem solving, guiding models through step-by-step solutions. ABench-Physics  (Zhang et al., 2025b) introduces highly challenging tasks to evaluate advanced capabilities, while PHYBench  (Qiu et al., 2025) targets symbolic computation. NEWTON  (Wang et al., 2023) assesses qualitative understanding, complementing the more quantitative approaches. Other benchmarks, such as LLM-SRBench  (Shojaee et al., 2025), focus specifically on quantitative scientific modeling. Together, these benchmarks show how models apply mathematical reasoning to well-defined physical laws, marking an important advance towards assessing real-world problem-solving. However, they remain largely confined to physics and have not been widely adopted for evaluating models in real-world engineering practices.
2.3 Benchmarks in Engineering
Current approaches to evaluating engineering reasoning are fractured, split between benchmarks that are overly broad and those that are deep but methodologically limited. On one hand, generalist benchmarks like MMLU  (Hendrycks et al., 2021a) and BIG-Bench  (Luo et al., 2024) include engineering topics, but their reliance on multiple-choice, knowledge-based questions primarily tests factual recall rather than the deep, procedural reasoning required for practical engineering problem-solving. On the other hand, more specialized engineering benchmarks, while deeper in scope, either remain confined to a single scientific discipline or lack a scalable, deterministic method for verifying the reasoning process itself. For example, EngiBench  (Felten et al., 2025) provides broad coverage but uses rubric-based scoring for its more complex problems, whereas FEA-Bench  (Li et al., 2025b) evaluates a model’s ability to use professional simulation software, yet this measures tool proficiency rather than fundamental reasoning.
The preceding review makes it clear that the field lacks a benchmark that captures the integrative, cross-disciplinary, and verifiable reasoning central to engineering. To address this critical gap, we introduce EngChain, a benchmark for verifiable, multi-step engineering reasoning. It uses a two-stage evaluation that combines quantitative verification of each intermediate step with a qualitative error diagnosis, covering 90 distinct problem types across nine domains and three engineering branches.
3 EngChain Taxonomy
EngChain covers a wide range of core engineering problems, focusing on domains that are central to real-world practice, which we show in Figure 2.
In Chemical Engineering, it targets the three pillars of chemical process design and analysis. Kinetics problems test the modeling of reaction rates and mechanisms, requiring reasoning about how systems evolve over time. Thermodynamics problems center on energy balances and phase behavior, capturing the constraints that govern feasibility and efficiency. Transport phenomena problems address momentum, heat, and mass transfer, linking microscopic principles to large-scale process behavior. In Electrical Engineering, it covers the full spectrum from fundamental physics to applied systems. At the foundation, it includes electromagnetics, which governs the behavior of fields and waves. Building on this, it addresses signals and systems, focusing on representation, transformation, and filtering of information. Finally, it extends to communications, where these principles are applied to the reliable transmission and processing of signals in real-world contexts. In Mechanical Engineering, it spans three core domains. Fluid mechanics problems address laminar and turbulent flow, testing how models capture momentum and energy transfer in fluids. Mechanics of materials emphasizes stress–strain relationships, deformation, and failure analysis, reflecting how structures respond to applied loads. Vibrations problems focus on natural frequencies and damping, capturing the dynamic stability of mechanical systems. Together, these areas test how forces, motion, and stability interact in physical systems.
Figure 2: The taxonomy of EngChain.
By integrating these principle subfields across the main branches of engineering, EngChain provides a comprehensive foundation of the analytical and procedural reasoning skills essential for engineering problem-solving.
4 EngChain Design and Methodology
To ensure the rigor, scalability, and long-term viability of the benchmark, we adopted a programmatic, template-based methodology. Unlike static datasets, which are finite and susceptible to being inadvertently included in the training corpora of future models, a programmatic approach offers several key advantages. It provides virtually limitless scalability and diversity, as a single template can generate thousands of unique problem instances. This design also ensures strong resistance to training data contamination, as it tests for generalizable reasoning on novel problems rather than the memorization of previously seen solutions.
4.1 Content Selection Framework
The selection of content for EngChain followed a principled, multi-stage methodology designed to ensure the benchmark is both representative of a formal engineering curriculum and grounded in foundational, real-world principles. The process involved a hierarchical selection of domains, areas, and finally, the specific problems to be templatized.
4.1.1 Domain and Area Selection
The nine core domains that form the backbone of EngChain were identified using a two-pronged approach. First, a thorough literature review was conducted, surveying established engineering taxonomies and curricular standards set by leading professional and accreditation bodies. This involved analyzing the program criteria from organizations like the Accreditation Board for Engineering and Technology (ABET, 2024), which mandates core competencies in areas like thermodynamics and circuit analysis for accredited degrees. We also reviewed the educational frameworks of major professional societies, such as the American Society of Mechanical Engineers (ASME, 2025), the Institute of Electrical and Electronics Engineers (IEEE, 2025), and the American Institute of Chemical Engineers (AIChE, 2025), to ensure our selected domains represent a consensus on the foundational knowledge required for professional practice.
Second, these findings were cross-referenced by querying an LLM (gemini-2.0-flash). The model was prompted to identify and rank the most fundamental, cornerstone domains within each of the three primary engineering branches. The full prompt used for this domain-level validation is provided in Appendix A. The final set of nine domains represents a strong consensus between traditional academic structures and the knowledge synthesized by modern AI. A similar two-pronged methodology was then used to identify the most critical areas within each domain: first, by extracting the core pedagogical units from the authoritative textbooks, and second, by cross-referencing this list using a LLM (gemini-2.0-flash). The full prompt used for this area-level validation is provided in Appendix B.
4.1.2 Template Selection
Once the domains and areas were defined, we developed the 90 problem templates. The core engineering principles, problem typologies, and typical constraints for these templates were sourced from the authoritative and widely-adopted textbooks in each field. This textbook-driven approach ensures that every problem in EngChain is pedagogically sound and central to engineering practice. A complete list of these foundational texts is provided in Appendix C.
Crucially, the number of templates developed for each area was not uniform. Instead, it was guided by a Pedagogical Significance Score assigned to each area. This score was programmatically determined by querying an expert-persona LLM (gemini-2.0-flash) to rate each area’s centrality to the curriculum on a 1-5 scale, as detailed in Appendix D. Foundational concepts with higher scores received more extensive coverage. For example, within the Reaction Kinetics domain, the area of Mole Balances (Significance Score: 5) was covered by five templates, while the more specialized application of Levenspiel Plot Interpretation (Significance Score: 1) was covered by a single template. Therefore, the varying number of templates per area is a direct and intentional outcome of our methodology. This ensures that EngChain’s structure reflects the hierarchical and foundational nature of engineering knowledge as it is taught and practiced.
@@ 4.2 Template Generation Pipeline
The template generation process follows a systematic pipeline as shown in Figure 3. First, we generate domain-aware parameters are generated. Next, the core engineering computations are performed. These results are then thoroughly validated to ensure they are physically plausible. Finally, the validated content is assembled into a natural language question and its corresponding step-by-step solution.
Figure 3: Our template generation pipeline.
We format the output strings to be both human-readable, like a textbook example, and machine-parsable for automated evaluation. After template functions are defined, we generate the problem instances by repeatedly calling each template, with every run using a unique random seed for reproducibility.
4.3 Template Structure and Outputs
The core of our benchmark is the template function, a standardized Python function that serves as the blueprint for a specific type of engineering problem. To ensure consistency, every template sticks to a simple, standardized signature i.e., it takes no arguments and returns a (question, solution) tuple. The question is a formatted string containing the problem statement to be presented to the LLM, whereas the solution is a formatted string containing the complete reasoning trace rather than only a numerical answer. A generic multi-disciplinary template structure is provided in Figure 4.
Figure 4: Generic multi-disciplinary template structure.
Each solution is structured like a textbook example such that the principles are stated, values are substituted, and intermediate steps are shown with consistent units.
4.4 Domain-Aware Parameterization and Sanity Checks
EngChain uses domain-aware parameterization, ensuring that problems are not just mathematically solvable but are also grounded in physical and engineering realism. This is done by constraining all generated values by the principles of the domain being tested.
For instance, in Chemical Engineering, problems use real reactants like Benzene, pre-validated stoichiometric equations, and temperature dependent heat capacity parameters rather than random constant values. In Electrical Engineering, parameters are grounded in fundamental constants (e.g., $c_{0}$, $\epsilon_{0}$) and real-world materials (e.g., Polyethylene), with signal properties procedurally generated to explicitly test concepts like aliasing. In Mechanical Engineering, problems co-select consistent properties for real materials (e.g., 6061-T6 Aluminum) and are context-aware, such as using Mercury for manometer problems. A comprehensive list of these parameterization examples is provided in Appendix E, and the authoritative data sources used are listed in Appendix F.
Furthermore, across all templates, these already constrained parameters are subjected to multi-layer sanity checks. For example, a chemical engineering template verifies that a final concentration is less than the initial one. A mechanical engineering problem ensures a calculated stress does not exceed the material’s yield strength. Similarly, an electrical engineering template confirms a wave’s propagation speed does not exceed the speed of light. These checks validate that the final combination of values results in a solvable and physically meaningful scenario.
4.5 AI-Assisted Quality Assurance
To ensure the correctness and scalability of the benchmark, we developed an AI-Assisted Quality Assurance workflow. While traditional unit tests are effective, they proved impractical for our generative framework due to the high number of templates and the complexity of verifying physically plausible outputs. Our quality assurance workflow automates this process by employing an LLM as a first-pass Peer Reviewer to validate new problem templates before their inclusion in the final benchmark.
The core of this workflow is a programmatic validation script that prepares a comprehensive validation package for the LLM. For each template, this package includes not only the template’s full Python source code but also three unique instances (question-solution pairs) generated by it. This context is crucial, as it allows the LLM to assess both the static logic of the code and the dynamic quality of its output. We then pass this package to the LLM via a carefully engineered prompt. As detailed in Appendix G, the prompt instructs the LLM to assume the persona of an expert engineering professor and to evaluate the template against a multi-axis rubric provided in Appendix H.
Upon receiving the results, our system programmatically parses and evaluates the scores. A template, $T$, is automatically approved if and only if the score assigned to it for every core category ($c_{i}$) in our rubric meets or exceeds a predefined threshold, $\theta$. We formalize this approval condition as:
$\forall c_{i}\in C,\text{Score}(T,c_{i})\geq\theta$
We established a high threshold of $\theta=4$ based on empirical testing against a control set of known-good and deliberately flawed templates (Examples are provided in Appendix L). Any template failing to meet this criterion is automatically flagged for human review. This threshold demonstrated an optimal balance between maximizing automation and reliably catching potential errors.
In practice, this automated review process proved highly effective. Out of the 90 total templates, only 5 (5.55%) were flagged by the LLM and required subsequent human review. [^1] This human review was performed by the authors, who possess relevant engineering domain expertise. While we did not calculate formal inter-rater reliability between the LLM and human reviewers for this quality assurance step, we manually inspected and corrected the flagged templates.
4.6 Symbolic Generation and Combinatorial Diversity
A core feature of EngChain is its programmatic generation, designed to create a problem space so vast that it effectively prevents evaluation through memorization. Rather than relying on a static set of problems, EngChain is built upon symbolic templates where parameters are systematically randomized within domain-aware, physical constraints.
This design provides strong resistance to training data contamination, as it makes it statistically improbable that an LLM has encountered any specific problem instance. Consequently, successful performance on EngChain is stronger evidence of a model’s ability to generalize its reasoning capabilities. The model is forced to solve each problem from first principles rather than relying on pattern matching or rote recall of seen solutions.

### v2

# EngTrace: A Symbolic Benchmark for Verifiable Process Supervision of Engineering Reasoning

@@ Abstract
Large Language Models (LLMs) are increasingly entering specialized, safety-critical engineering workflows governed by strict quantitative standards and immutable physical laws, making rigorous evaluation of their reasoning capabilities imperative. However, existing benchmarks such as MMLU, MATH, and HumanEval assess isolated cognitive skills, failing to capture the physically grounded reasoning central to engineering, where scientific principles, quantitative modeling, and practical constraints must converge. To enable verifiable process supervision in engineering, we introduce EngTrace, a symbolic benchmark comprising 90 templates across three major engineering branches, nine core domains and 20 distinct areas. Through domain-aware parameterization, we generate 1,350 unique, contamination-resistant test cases to stress-test generalization. Moving beyond outcome matching, we introduce a verifiable two-stage evaluation framework that uses a tiered protocol to validate intermediate reasoning traces alongside final answers through automated procedural checks and a heterogeneous AI Tribunal. Our evaluation of 24 leading LLMs reveals a distinct trade-off between numeric precision and trace fidelity, identifying a complexity cliff where abstract mathematical pre-training fails to translate into the integrative reasoning required for advanced engineering tasks.
EngTrace: A Symbolic Benchmark
for Verifiable Process Supervision of Engineering Reasoning
Ayesha Gull1 Muhammad Usman Safder1 Rania Hossam2 Fan Zhang3 Veselin Stoyanov2 Preslav Nakov2 Zhuohan Xie2 1Namal University 2MBZUAI 3The University of Tokyo Project Code App
@@ 1 Introduction
With the expansion of LLMs into high-stakes domains such as engineering, the need for their comprehensive evaluation has become paramount (Zhao et al., 2023; Xie et al., 2023a). Progress is currently gated by benchmark quality; without tests that enforce physically grounded reasoning, distinguishing genuine capability from sophisticated mimicry remains a challenge.
@@
Figure 1: An example from EngTrace demonstrating CSTR Volume Calculation. A symbolic template generates a unique problem instance paired with a gold-standard reasoning trace to enable verifiable process supervision.
Extending the symbolic-template paradigm pioneered in mathematics (Mirzadeh et al., 2024) and recently in finance (Xie et al., 2025) to the interdependent physical systems and specialized design constraints of engineering, we introduce EngTrace, a new benchmark designed to enable verifiable process supervision of reasoning in engineering problem-solving. It is built on a foundation of symbolic templates, each capable of programmatically generating a vast number of unique problem instances. Figure 1 illustrates a Continuous Stirred-Tank Reactor (CSTR) volume calculation example from our benchmark. Unlike static benchmarks, this generated instance requires synthesizing interdependent stoichiometry and reaction kinetics variables to derive the reactor size required for a specific outlet flow rate, yielding a complex reasoning trace fully verifiable against the template’s ground truth.
EngTrace addresses two critical gaps that compromise the validity of current LLM evaluation paradigms for engineering: benchmark saturation and disciplinary fragmentation. First, static benchmarks like Glue (Wang et al., 2019) and BBH (Kazemi et al., 2025) are increasingly prone to model saturation (Ott et al., 2022) and data contamination (Deng et al., 2024). EngTrace addresses these risks through symbolic templates and domain-aware parameterization, generating unique, physically grounded problems that prevent rote memorization. Second, most benchmarks evaluate critical skills in disciplinary silos. General knowledge benchmarks (e.g., MMLU (Hendrycks et al., 2020)) test broad factual recall, while specialized benchmarks (e.g., Math (Hendrycks et al., 2021b) or HumanEval (Chen et al., 2021)) assess abstract logic and algorithmic translation, respectively. This fragmentation is limiting for engineering which is fundamentally an integrative discipline that requires the synthesis of these isolated skills, combining scientific principles, mathematical modeling, and practical constraints (e.g., Moaveni, 2019; Dym et al., 2012). EngTrace addresses this structural limitation through integrative reasoning templates. Unlike existing benchmarks (e.g., Zhou et al., 2025; Mudur et al., 2025) that rely solely on outcome matching, our templates demand holistic, procedural reasoning. This is a crucial requirement in engineering, where a flawed process can lead to catastrophic failure (Leveson, 2012).
In summary, our contributions are threefold: (1) We introduce EngTrace, a gold-standard symbolic benchmark for verifiable process supervision, comprising 90 problem templates spanning three major engineering branches, nine core domains, and 20 distinct areas. (2) We propose a verifiable two-stage evaluation framework that employs a tiered verification protocol utilizing automated procedural checks and a multi-model AI tribunal to assess reasoning process quality in addition to final answer accuracy. (3) We evaluate 24 models across various categories, revealing critical weaknesses in reasoning quality, with non-frontier models failing to generalize beyond basic patterns and performing poorly under complex engineering constraints.
@@ 2 Related Work
Benchmarks in Mathematics and Coding.
Extensive work has established baselines for isolated cognitive skills. In mathematics, benchmarks ranging from GSM8K (Cobbe et al., 2021) and Math (Hendrycks et al., 2021b) to HardMath (Fan et al., 2024), Putnam-Axiom (Gulati et al., 2024), and GSM-Symbolic (Mirzadeh et al., 2024) rigorously test abstract logical deduction and reasoning stability. Similarly, coding evaluations span from function-level generation in HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) to real-world software engineering in SWE-Bench (Jimenez et al., 2024). While foundational, these benchmarks lack engineering utility, as proficiency in abstract logic does not translate to reliable reasoning under physical constraints (Heesch et al., 2025). Recently, FinChain (Xie et al., 2025) utilized symbolic templates to evaluate financial reasoning. While this demonstrated the utility of verifiable workflows, EngTrace extends the approach to physically grounded simulation, where reasoning traces are governed by immutable conservation laws and interdependent system dynamics.
Benchmarks in the Physical Sciences.
To bridge the gap between abstract mathematics and physical reality, recent benchmarks have pivoted to the physical sciences. While UGPhysics (Xu et al., 2025) and PhysReason (Zhang et al., 2025a) effectively test the application of reasoning to natural laws, others target advanced symbolic (Zhang et al., 2025b; Qiu et al., 2025) or qualitative (Wang et al., 2023; Shojaee et al., 2025) understanding. Despite this progress, these benchmarks remain confined to theoretical physics and lack the integrative context of engineering.
Benchmarks in Engineering.
This fragmentation extends to engineering evaluation. While generalist suites like MMLU (Hendrycks et al., 2021a),BIG-Bench (Luo et al., 2024), and SuperGPQA (Du et al., 2025) include engineering topics, their reliance on knowledge-based multiple-choice questions restricts assessment to factual recall rather than procedural fidelity. Conversely, specialized benchmarks offer depth but are constrained by narrow disciplinary scopes or a lack of scalable reasoning verification. For instance, specialized benchmarks like TransportBench, APBench, and EEE-Bench are siloed within narrow sub-disciplines (Syed et al., 2024; Chen et al., 2025; Cheng et al., 2024; Li et al., 2025; Liu and others, 2025). Others focus on tangential skills, such as design generation (Guo et al., 2025b) or software proficiency (Mudur et al., 2025), rather than fundamental reasoning, while broader efforts like EngiBench (Zhou et al., 2025) rely on subjective rubric-based scoring.
Evidently, existing benchmarks remain limited to outcome matching and fail to validate the physically grounded reasoning required for engineering. EngTrace addresses this limitation by introducing a symbolic benchmark that pairs unique problem instances with gold-standard reasoning traces to enable verifiable process supervision.
3 EngTrace Design and Methodology
3.1 Taxonomy and Content Selection
The taxonomy of EngTrace, illustrated in Figure 2, systematically covers nine core domains and 20 distinct areas across three engineering branches. This structure is not arbitrarily defined but follows a principled, two-pronged methodology to ensure it mirrors formal engineering curricula.
Figure 2: EngTrace’s Hierarchical Taxonomy spanning nine domains across three branches: process equilibrium and dynamics (Chemical), field theory and signal transmission (Electrical), and physical mechanics and material properties (Mechanical).
Domain and Area Selection.
To identify core domains and critical areas, we synthesized engineering curricular standards from ABET (ABET, 2024) and major professional societies (ASME (ASME, 2025), IEEE (IEEE, 2025), AIChE (AIChE, 2025)). We validated these foundational units by cross-referencing them with an expert-persona LLM to ensure alignment between traditional pedagogy and modern AI capabilities (prompts provided in Appendix A and Appendix B).
Template Selection.
We developed 90 symbolic templates by sourcing core engineering principles, problem typologies, and typical constraints from foundational textbooks (e.g., Fogler, 2020; Ulaby and Ravaioli, 2015), ensuring pedagogical grounding (see Appendix C for the complete list of source texts). The template distribution per area is non-uniform and guided by a Pedagogical Significance Score (1–5) derived through an expert-persona LLM (detailed prompt provided in Appendix D). This weighting ensures hierarchical fidelity: for instance, within the Reaction Kinetics domain, the core area of Mole Balances (Score: 5) includes five templates, while the specialized Levenspiel Plot Interpretation (Score: 1) is represented by a single template.
@@ 3.2 Template Generation Pipeline
The core of EngTrace is the symbolic template, a Python blueprint for generating unique, physically grounded engineering problems. As shown in Figure 3, the pipeline first samples domain-aware parameters from authoritative data sources (Appendix E) and fundamental principles (see detailed examples in Appendix F). It then performs core computations with multi-layer sanity checks to derive computationally verified ground-truth traces. These are serialized into natural language (question, solution) pairs, creating a dynamic problem space that effectively resists data contamination.
3.3 Template Validation
To establish EngTrace as a high-fidelity “Gold Standard” benchmark, we utilize a robust two-stage validation process that pairs automated heuristic auditing with expert manual verification.
Stage 1: Automated AI Pre-Screening.
Building upon the “Panel of LLM” (PoLL) methodology (Verga et al., 2024), we construct an “AI Tribunal” comprising three frontier reasoning models: GPT-5 (OpenAI, 2025a), Claude Opus 4.5 (Anthropic, 2025b), and Gemini 3 (Google, 2025). Each model acts as an independent judge, utilizing the specific prompt provided in Appendix G, to evaluate the source code and generated instances against the strict multi-axis rubric detailed in Appendix H.
To mitigate potential hallucinations and ensure output reliability, we apply a hybrid aggregation strategy requiring each template $T$ to satisfy three consensus conditions. First, the Median Quality Score ($S_{med}$) across the tribunal must reach a threshold $\theta=4$ for every evaluative dimension $d\in D$:
$\forall d\in D,\quad S_{med}^{(d)}=\text{median}(S_{m_{1}}^{(d)},S_{m_{2}}^{(d)},S_{m_{3}}^{(d)})\geq 4$ (1)
Second, we enforce a Majority Vote safety check, discarding any template where more than one judge identifies a violation ($\sum F_{i}\geq 2$, where $F\in\{0,1\}$).
Finally, we quantify uncertainty using a Disagreement Score ($\sigma_{max}$), defined as the maximum standard deviation across all dimensions. Templates exhibiting high variance are flagged for manual review to resolve potential logical ambiguities:
$\sigma_{max}=\max_{d\in D}\left(\sqrt{\frac{1}{3}\sum_{i=1}^{3}(S_{m_{i}}^{(d)}-\mu^{(d)})^{2}}\right)\leq 0.5$ (2)
where $\mu^{(d)}$ represents the mean score across judges for dimension $d$.
Stage 2: Expert Human Certification.
Templates passing the AI tribunal undergo a double-blind review by domain experts through a custom Streamlit interface presented in Appendix I. We assess Inter-Annotator Agreement per branch using Fleiss’ Kappa ($\kappa$) for binary admissibility (Approve/Reject) and Krippendorff’s Alpha ($\alpha$) for ordinal quality ratings (1–5), the latter accounting for disagreement magnitude and missing data.
AI-Human Alignment.
To assess validation reliability, we measure correspondence between the AI Tribunal and human experts using False Positive Rate (FPR), to quantify filter safety (templates approved by AI but rejected by human experts), and Spearman’s Rank Correlation ($\rho$) to evaluate monotonic ranking consistency, chosen over Pearson to better handle the ordinal nature and high-quality skew of the data. Detailed results are provided in Appendix J.
Figure 3: The EngTrace Template Generation Pipeline maps domain-aware constraints to serialized question-solution pairs via physically-validated computation, illustrated here using a CSTR volume calculation example.